# On a group of normalized solutions of the higher-dimensional Prouhet–Tarry–Escott problem

### v1

@@ 1. Introduction
Alpers–Tijdeman [Alpers-Tijdeman] established a geometric framework for $\operatorname{PTE}_{2}$ including the explicit construction of ideal solutions. This framework was generalized to $\operatorname{PTE}_{\geq 3}$ by Ghiglione [Ghiglione] via geometric techniques related to discrete tomography. In [MS2025+], the second and third authors constructed a family of rational ideal solutions from a certain combinatorial object called ellipsoidal design; for the definition of ellipsoidal designs, see Definition 3.1.
@@
Suppose that $A]=_{m}^{n}[B]$ is a (not necessarily disjoint) solution of $\operatorname{PTE}_{r}$ and $\operatorname{rank}A=\operatorname{rank}B=r$, where AA and $B$ are regarded as the $r\times n$ matrices whose columns are $(a_{i1},\ldots,a_{ir})^{T}$ and $(b_{i1},\ldots,b_{ir})^{T}$ respectively. Then, $(A,B)$ is called a normalized solution if the $r\times 2n$ matrix [AB][A\;B] concatenating $A$ and $B$ is in the reduced row echelon form.
@@
The full rank condition of $A$ and $B$ in Definition 1.4 is natural because most of the solutions including those of $\operatorname{PTE}_{2}$ in [Alpers-Tijdeman, MS2025+] satisfy this condition. Moreover, Alpers–Tijdeman [Alpers-Tijdeman, Theorem 8] established a construction method of solutions of $\operatorname{PTE}_{2}$ from points which have equal $X$-rays along with different directions. For $(p,q)\in\mathbb{Z}^{2}$ and $B\subset\mathbb{Z}^{2}, AA$ and $B$ have equal $X$-rays along with $\operatorname{lin}\{(p,q)\}$, where $\operatorname{lin}\{(p,q)\}=\{(\lambda p,\lambda q)\mid\lambda\in\mathbb{R}\}$, if they have an equal number of points on each line parallel to $\operatorname{lin}\{(p,q)\}$. This construction was generalized to $\operatorname{PTE}_{r}$ by Ghiglione [Ghiglione, Theorems 4.3.17 and 4.3.20].
@@
(2) $Q=\operatorname{diag}\left[1,3,\cdots,\frac{n(n-1)}{2}\right].$
@@ 2. Proof of the main theorem
Let $m,n\geq 2$. Then it holds that
$N_{2}(n)=\{A\in M_{n}(\mathbb{Q})\mid AA^{T}=I_{n},\;A\mathbf{1}=\mathbf{1}\}.$
@@
Clearly, the left hand side contains the right hand side. Suppose $C\in N_{2}(n)$. Then there exist $A,B\in M_{2}(\mathbb{Q})$ such that $C=A^{-1}B$ and $A]=_{2}^{n}[B]$. Since $A]=_{2}^{n}[B]$, we have $AA^{T}=BB^{T}$ and $A\mathbf{1}=B\mathbf{1}$. Therefore, we have $(A^{-1}B)(A^{-1}B)^{T}=I_{n}$ and $A^{-1}B\mathbf{1}=\mathbf{1}$. By the representation of the right hand side, $N_{2}(n)$ forms a group under matrix multiplication. ∎
Next, we prove the key lemma.
Lemma 2.2.
Let $n\geq 2$ be an integer and $A\in N_{2}(n)$. Let $P\in\operatorname{GL}_{n}(\mathbb{Q})$ be a matrix whose first column is $\mathbf{1}$, and whose any two distinct columns are mutually orthogonal. Then there exists $B^{\prime}\in M_{n-1}(\mathbb{Q})$ such that
(3) $P^{-1}AP=\begin{bmatrix}1&\mathbf{0}^{T}\\ \mathbf{0}&B^{\prime}\end{bmatrix}$.
Proof.
Write
$P=\begin{bmatrix}\mathbf{1}&\mathbf{b}_{1}&\cdots&\mathbf{b}_{n-1}\end{bmatrix}$,
where any two distinct columns of $P$ are mutually orthogonal. Then, since $A\mathbf{1}=\mathbf{1}$ by Proposition 2.1 and $P\mathbf{e}_{1}=\mathbf{1}$, we have
$P^{-1}AP=\begin{bmatrix}\mathbf{e}_{1}&P^{-1}A\mathbf{b}_{1}&\cdots&P^{-1}A\mathbf{b}_{n-1}\end{bmatrix}$.
Thus, it is sufficient to prove that $P^{-1}A\mathbf{b}_{k}\in\langle\mathbf{e}_{1}\rangle^{\perp}=\langle\mathbf{e}_{2},\ldots,\mathbf{e}_{n}\rangle$ for all $1\leq k\leq n-1$.
First, note that
$A\langle\mathbf{1}\rangle^{\perp}\subset\langle\mathbf{1}\rangle^{\perp}.$
Indeed, since $A\mathbf{1}=\mathbf{1}$ and $AA^{T}=I_{n}$ by Proposition 2.1, we have
for all $\mathbf{x}\in\langle\mathbf{1}\rangle^{\perp}$. Since
$A\mathbf{b}_{k}\in\langle 1\rangle^{\perp}=\langle\mathbf{b}_{1},\ldots,\mathbf{b}_{n-1}\rangle$
for all $1\leq k\leq n-1$, we obtain
$P^{-1}A\mathbf{b}_{k}\in\langle P^{-1}A\mathbf{b}_{1},\ldots,P^{-1}A\mathbf{b}_{n-1}\rangle\subset\langle P^{-1}\mathbf{b}_{1},\ldots P^{-1}\mathbf{b}_{n-1}\rangle=\langle\mathbf{e}_{2},\ldots,\mathbf{e}_{n}\rangle.$
This completes the proof. ∎
@@
Let $\mathbf{b}_{k}=[1,\ldots,1,-k,0,\ldots,0]^{T}\in M_{n,1}(\mathbb{Q})$, whose $(k+1)$-st component is k-k, and
$P:=\begin{bmatrix}\mathbf{1}&\mathbf{b}_{1}&\cdots&\mathbf{b}_{n-1}\end{bmatrix}$.
Then $P\in\operatorname{GL}_{n}(\mathbb{Q})$ and its any two distinct columns are mutually orthogonal.
Let $A\in N_{2}(n)$. Then by Lemma 2.2, there exists $B^{\prime}\in M_{n-1}(\mathbb{Q})$ such that
$P^{-1}AP=\begin{bmatrix}1&\mathbf{0}^{T}\\ \mathbf{0}&B^{\prime}\end{bmatrix}$.
Since $N_{2}(n)\simeq P^{-1}N_{2}(n)P$, it is sufficient to prove that
$P^{-1}N_{2}(n)P=\left\{\begin{bmatrix}1&\mathbf{0}^{T}\\ \mathbf{0}&B^{\prime}\end{bmatrix}\in M_{n}(\mathbb{Q})\middle|(B^{\prime})^{T}Q^{\prime}B^{\prime}=Q^{\prime}\right\}=:G$
with
(4) $Q^{\prime}:=\begin{bmatrix}(\mathbf{b}_{i},\mathbf{b}_{j})\end{bmatrix}=\operatorname{diag}[|\mathbf{b}_{1}|^{2},\cdots,|\mathbf{b}_{n-1}|^{2}]=\operatorname{diag}[2,6,\cdots,n(n-1)].$
Indeed, let $A\in N_{2}(n)$ and $B:=P^{-1}AP$. Then we have
[…]
Therefore, we have $(B^{\prime})^{T}Q^{\prime}B^{\prime}=Q^{\prime}$, i.e., $P^{-1}AP\in G$.
Next, take
$B=\begin{bmatrix}1&\mathbf{0}^{T}\\ \mathbf{0}&B^{\prime}\end{bmatrix}\in G.$
Then we have
$(PBP^{-1})(PBP^{-1})^{T}$ $=PB(P^{T}P)^{-1}B^{T}P^{T}$
$=PB\begin{bmatrix}n&\mathbf{0}^{T}\\ \mathbf{0}&Q^{\prime}\end{bmatrix}^{-1}B^{T}P^{T}$
$=P\begin{bmatrix}\frac{1}{n}&\mathbf{0}^{T}\\ \mathbf{0}&B^{\prime}Q^{\prime-1}(B^{\prime})^{T}\end{bmatrix}P^{T}$
$=P\begin{bmatrix}\frac{1}{n}&\mathbf{0}^{T}\\ \mathbf{0}&(Q^{\prime})^{-1}\end{bmatrix}P^{T}$
$=P(P^{T}P)^{-1}P^{T}$
$=I_{n}.$
Since $B^{\prime}\in O_{n-1}(Q^{\prime},\mathbb{Q})$, the fourth equality follows from
$B^{\prime}Q^{\prime-1}(B^{\prime})^{T}=B^{\prime}((B^{\prime})^{T}Q^{\prime}B^{\prime})^{-1}(B^{\prime})^{T}=(Q^{\prime})^{-1}.$
The second and fifth equalities follow from the last equality of (5).
By normalization, $Q^{\prime}$ and
$Q:=\frac{1}{2}Q^{\prime}=\operatorname{diag}\left[1,3,\cdots,\frac{n(n-1)}{2}\right]$
are similar (Definition 2.3). This completes the proof. ∎
@@ 3. Further remarks and future works
The tightness was first defined by Delsarte et al. [DGS1977] for spherical designs. Since an ellipsoid and a sphere are homeomorphic, we can similarly define the tightness for ellipsoidal designs in general. However, by considering rational points, we can find an essential gap between spheres and ellipses. Namely, there exists a rational tight design in $\mathbb{R}^{2}$ only if $(x,y)Q(x,y)^{T}=x^{2}+y^{2}$ or $x^{2}+xy+y^{2}$ (see [MS2025+, Theorem 4.4]). It is still an open question whether there exists a rational (not necessarily tight) $4$-design on $\mathbb{S}^{1}$, whereas there exist infinitely many rational tight 55-designs on the ellipse $x^{2}+xy+y^{2}=1$. In [MS2025+, Theorem 5.5], the second and third authors established a construction method of rational ideal solutions of $\operatorname{PTE}_{2}$ by utilizing rational tight designs over the ellipse $x^{2}+xy+y^{2}=1$. Since $x^{2}+xy+y^{2}=(x+y/2)^{2}+3(y/2)^{2}$, the two quadratic forms $x^{2}+xy+y^{2}$ and $x^{2}+3y^{2}$ are equivalent. The latter quadratic form appears in the case $n=3$ of Theorem 1.6.

### v2

@@ 1. Introduction
Alpers–Tijdeman [Alpers-Tijdeman] established a geometric framework for $\operatorname{PTE}_{2}$ including the explicit construction of ideal solutions. This framework was generalized to $\operatorname{PTE}_{\geq 3}$ by Ghiglione [Ghiglione] via geometric techniques related to discrete tomography and techniques involving Gröbner bases. Matsumura and Sawa [MS2025+] constructed a family of rational ideal solutions from a certain combinatorial object called ellipsoidal design; for the definition of ellipsoidal designs, see Definition 3.1.
@@
Suppose that $A]=_{m}^{n}[B]$ is a (not necessarily disjoint) solution of $\operatorname{PTE}_{r}$ and $\operatorname{rank}A=\operatorname{rank}B=r$, where $A$ and $B$ are regarded as the $r\times n$ matrices whose columns are $(a_{i1},\ldots,a_{ir})^{T}$ and $(b_{i1},\ldots,b_{ir})^{T}$ respectively. Then, $(A,B)$ is called a normalized solution if the $r\times 2n$ matrix [AB][A\;B] concatenating $A$ and $B$ is in the reduced row echelon form.
@@
The full rank condition of $A$ and $B$ in Definition 1.4 is natural because most of the solutions including those of $\operatorname{PTE}_{2}$ in [Alpers-Tijdeman, MS2025+] satisfy this condition. In particular, Alpers–Tijdeman [Alpers-Tijdeman, Theorem 8] established a construction method of solutions of $\operatorname{PTE}_{2}$ from points which have equal $X$-rays along with different directions. For $(p,q)\in\mathbb{Z}^{2}$ and $B\subset\mathbb{Z}^{2}, AA$ and $B$ have equal $X$-rays along with $\operatorname{lin}\{(p,q)\}$, where $\operatorname{lin}\{(p,q)\}=\{(\lambda p,\lambda q)\mid\lambda\in\mathbb{R}\}$, if they have an equal number of points on each line parallel to $\operatorname{lin}\{(p,q)\}$. This construction was generalized to $\operatorname{PTE}_{r}$ by Ghiglione [Ghiglione, Theorems 4.3.17 and 4.3.20].
@@
(2) $Q=\operatorname{diag}\left[1,3,\ldots,\frac{n(n-1)}{2}\right].$
@@ 2. Proof of the main theorem
Let $n\geq 2$. Then it holds that
$N_{2}(n)=\{A\in O_{n}(\mathbb{Q})\mid A\mathbf{1}=\mathbf{1}\}.$
@@
Let $A$ be in the right hand side. Since $A\mathbf{1}=\mathbf{1}$ and $AA^{T}=I_{n}$ imply the conditions of $PTE_{n}$ of degree $1$ and $2$ respectively, it follows that $A=I_{n}^{-1}A\in N_{2}(n)$.
Suppose $C\in N_{2}(n)$. Then there exist $A,B\in M_{2}(\mathbb{Q})$ such that $C=A^{-1}B$ and $A]=_{2}^{n}[B]$. Since $A]=_{2}^{n}[B]$, we have $AA^{T}=BB^{T}$ and $A\mathbf{1}=B\mathbf{1}$. Therefore, we have $(A^{-1}B)(A^{-1}B)^{T}=I_{n}$ and $A^{-1}B\mathbf{1}=\mathbf{1}$. By the representation of the right hand side, $N_{2}(n)$ forms a group under matrix multiplication. ∎
@@
By Proposition 2.1, we have
$N_{2}(n)=\{A\in O_{n}(\mathbb{Q})\mid A\mathbf{1}=\mathbf{1}\}.$
Let $A\in N_{2}(n)$ . Then for all $\mathbf{x}\in\langle\mathbf{1}\rangle^{\perp}$, we have
since $A\mathbf{1}=\mathbf{1}$ and $AA^{T}=I_{n}$. Therefore, $A$ defines an orthogonal transform $\phi:\langle\mathbf{1}\rangle^{\perp}\to\langle\mathbf{1}\rangle^{\perp}$ that preserves the rationality. Here, $\langle\mathbf{1}\rangle$ is the $\mathbb{Q}$-vector space spanned by $\mathbf{1}$ and $\langle\mathbf{1}\rangle^{\perp}$ is its orthogonal complement.
Let $\mathbf{e}_{k}=[1,\ldots,1,-k,0,\ldots,0]^{T}\in M_{n,1}(\mathbb{Q})$, whose $(k+1)$-st component is k-k. Then $\{\mathbf{e}_{k}\}_{k=1}^{n-1}$ forms a basis of $\langle\mathbf{1}\rangle^{\perp}$. Let $B$ be the representation matrix of $\phi$ with respect to the basis $\{\mathbf{e}_{k}\}_{k=1}^{n-1}$ and
$2Q:=[(\mathbf{e}_{i},\mathbf{e}_{j})_{ij}]=2\operatorname{diag}\left[1,3,\ldots,\frac{n(n-1)}{2}\right].$
If we denote $\mathbf{x}=\xi_{1}\mathbf{e}_{1}+\dots+\xi_{n-1}\mathbf{e}_{n-1}\in\langle\mathbf{1}\rangle^{\perp}$ and $\boldsymbol{\xi}:=(\xi_{1},\ldots,\xi_{n-1})$, then we have
$\boldsymbol{\xi}^{T}2Q\boldsymbol{\xi}=(\mathbf{x},\mathbf{x})=(A\mathbf{x},A\mathbf{x})=(B\boldsymbol{\xi})^{T}2Q(B\boldsymbol{\xi}).$
Thus, we obtain $B^{T}QB=Q$. Therefore, we have an isomorphism
$N_{2}(n)\simeq O_{n-1}(Q,\mathbb{Q}),\;A\mapsto B.$
Remark 2.2.
Theorem 1.6 can be proven without explicit computation, except for a specific orthonormal basis $\{\mathbf{e}_{k}\}_{k=1}^{n-1}$. Indeed, $\phi$ is equivalent to an isometry on $\langle\mathbf{1}\rangle^{\perp}$, i.e., it preserves the quadratic form $2Q$, and the group of isometries are isomorphic to $O_{n-1}(2Q)=O_{n-1}(Q)$. See [Gerstein2008, Definition 2.8 and p. 26] for details.
@@ 3. Further remarks and future works
The tightness was first defined by Delsarte et al. [DGS1977] for spherical designs. Since an ellipsoid and a sphere are homeomorphic, we can similarly define the tightness for ellipsoidal designs in general. However, by considering rational points, we can find an essential gap between spheres and ellipses. Namely, there exists a rational tight design in $\mathbb{R}^{2}$ only if $(x,y)Q(x,y)^{T}=x^{2}+y^{2}$ or $x^{2}+xy+y^{2}$ (see [MS2025+, Theorem 4.4]). It is still an open question whether there exists a rational (not necessarily tight) $4$-design on $\mathbb{S}^{1}$, whereas there exist infinitely many rational tight $5$-designs on the ellipse $x^{2}+xy+y^{2}=1$. Matsumura and Sawa [MS2025+, Theorem 5.5] established a construction method of rational ideal solutions of $\operatorname{PTE}_{2}$ by utilizing rational tight designs over the ellipse $x^{2}+xy+y^{2}=1$. Since $x^{2}+xy+y^{2}=(x+y/2)^{2}+3(y/2)^{2}$, the two quadratic forms $x^{2}+xy+y^{2}$ and $x^{2}+3y^{2}$ are equivalent. The latter quadratic form appears in the case $n=3$ of Theorem 1.6.
@@
Acknowledgements. The authors thank Hiroki Shimakura for helpful comments on the earlier draft.